# Examining Algorithmic Curation on Social Media: An Empirical Audit of Reddit’s r/popular Feed

### v1

@@ 9 Conclusion
Now that algorithmic curation has become integral to online ecosystems, examining it and its effects on user behavior becomes critical. In this paper, we conducted a comprehensive empirical audit of one such system: Reddit’s r/popular feed. Through this, we successfully quantified how recent commenting rate and other factors influence algorithmic ranking on r/popular, as well as how rank/position on r/popular correlates with subsequent engagement and undesirable behavior. Our findings are based on millions of snapshots of the top 100 ranks on r/popular consistently collected over 11 months, an approach that can be applied to other platforms. Additionally, we discussed the implications of our findings for stakeholders, including content creators, highlighted moderators’ lack of agency in community curation systems, and proposed future research directions on algorithmic curation amid reduced data access. All in all, studying algorithmic curation goes back to user agency: should users have control over how they interact with content and how their content is distributed online? To address this, we must first understand how these systems are embedded in our online environments, thus enabling us to make informed decisions about governing them.

### v3

@@ 9 Conclusion
Now that algorithmic curation has become integral to online ecosystems, examining it and its effects on user behavior becomes critical. In this paper, we conducted a comprehensive empirical audit of one such system: Reddit’s r/popular feed. Through this, we successfully quantified how recent commenting rate and other factors influence algorithmic ranking on r/popular, as well as how rank/position on r/popular impact subsequent engagement and undesirable behavior. Our findings are based on millions of snapshots of the top 100 ranks on r/popular consistently collected over 11 months, an approach that can be applied to other platforms. Additionally, we discussed the implications of our findings for stakeholders, including content creators, highlighted moderators’ lack of agency in community curation systems, and proposed future research directions on algorithmic curation amid reduced data access. All in all, studying algorithmic curation goes back to user agency: should users have control over how they interact with content and how their content is distributed online? To address this, we must first understand how these systems are embedded in our online environments, thus enabling us to make informed decisions about governing them.
@@
Appendix B Stratified Propensity Score Matching
| Strata | Samples | ||||||
| RQ | Variable | Strength | Cutoff (Control, Treatment) | Pre-Filter | Post | Per Stratum | Total |
| 1 & 2 | Rec. Comments | Weak | (11), | 200 | 97 | 7,333 | 711,297 |
| 1 & 2 | Rec. Comments | Strong | (11), (24) | 200 | 14 | 5,679 | 79,506 |
| 1 & 2 | Rec. Und. | Weak | 0, 1 | 200 | 124 | 7,348 | 911,174 |
| 1 & 2 | Rec. Und. | Strong | 0, (5) | 200 | 39 | 2,283 | 153,342 |
| 1 & 2 | Prop. Und. | Weak | (18.75%), | 200 | 166 | 6,989 | 1,160,062 |
| 1 & 2 | Prop. Und. | Strong | (18.75%), (31.25%) | 200 | 89 | 5,260 | 468,167 |
| 3 | Rank Movement | Weak | 0 (down), 0 (up) | 100 | 93 | 2,521 | 234,510 |
| 3 | Rank Movement | Strong | 0 (down), (up 3) | 100 | 87 | 1,672 | 147,283 |
Appendix C GLMs/Hierarchical Bayesian Models
Here, we detail the models used for our causal inference analyses in Sections 5-7.
RQ1
Recall that we tested multiple treatments and outcomes, with four treatments (TC-W, TC-S, TU-W, TU-S), and three outcomes (top 10, 20, and 50). Each treatment/outcome pair was modeled separately for a total of separate models.
For a particular treatment, let if the -th observation was in the treatment condition, and if it was in the control condition. Then, for a particular outcome , the probability that a post would appear in the top 10 (20, or 50) in the next snapshot was modeled as follows:
Here, denotes the stratum that the -th observation was assigned to, and with separate intercepts and coefficients for each stratum. Each intercept represents the baseline (i.e., control) log-odds of appearing in the top 10 (20, or 50) within the -th stratum. Each coefficient corresponds to the treatment effect in the -th stratum. To compute the average treatment effect across all strata, the coefficients were sampled from a hyper-distribution with mean and standard deviation (the intercepts were pooled in a similar manner):
The scalar thus represents the average treatment effect. Weakly informative priors were placed on these hyper-parameters (with similar priors on the hyper-parameters for the intercepts):
A posterior distribution for the average treatment effect was sampled for each treatment/outcome pair with Monte-Carlo Markov Chain sampling using PyMC. 333https://github.com/pymc-devs/pymc
RQ2
Again, recall that we tested multiple conditions and outcomes, with the same four treatment as before, and three outcomes (movement up, movement down, no movement). Each treatment was modeled separately, but the outcomes were modeled jointly using multinomial logistic regressions, yielding a total of just 4 models, one for each treatment.
Let the possible values of the outcome be up, none, and down. For each observation , let be the odds that the post would move up the feed in the next snapshot against the not moving in the next snapshot, and similarly for . We modeled these odds as follows:
Again, denotes the stratum that the -th observation was assigned to, and with separate intercepts ( and ) and coefficients (, ) for each stratum. Each intercept represents the baseline (i.e., control) log-odds of moving up the feed against no movement within the -th stratum. Each coefficient corresponds to the treatment effect in the -th stratum. The similar applies to and . The average treatment effect was modeled in a similar way as above, with similar priors on the hyper-parameters. Finally, posterior distributions for average treatment effects and were sampled for each treatment.
RQ3
For RQ3, we combined consecutive snapshots where there was no movement in rank into a single observation, yielding observations that spanned varying amounts of time, and which had rank movements just prior and just after the spanned time window. For the first two parts, where we investigated the impact on the number of overall comments and the number of undesired comments, we followed a similar procedure: Let be the time spanned during the -th observation window, and let be the number of comments received during that window. We modeled using a negative binomial distribution with dispersion parameter :
Here, can be interpreted as the predicted commenting rate for the -th observation, which is multiplied by to get the expected number of comments over that time window. Again, denotes the stratum that the -th observation was assigned to, and with separate intercepts and coefficients for each stratum. Each intercept represents the baseline (i.e. control) log-odds of commenting rate within the -th stratum. Each coefficient corresponds to the treatment effect in the -th stratum. The average treatment effect was modeled in a similar way as above, with similar priors on the hyper-parameters. Finally, posterior distributions for average treatment effect were sampled for each treatment.
For the third part, where we investigated the impact on the proportion of undesired comments, we followed a slightly different procedure. Let be the number of undesired comments and let be the total number of comments received in the -th observation window. We modeled as follows:
Here, can be interpreted as the predicted proportion of undesired comments out of all comments received in the -th observation. Each intercept represents the baseline log ratio of undesired comments, and each coefficient corresponds to the treatment effect. Average treatment effects for each treatment were computed in a similar manner as before.